Shift report 20011022 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Oct 22 2001 - 00:03:57 EDT

  • Next message: Flemming Videbaek: "Re: Shift report 20011021 08:00-16:00 - pii6,pii7"

    ------------------------------------------------------------
    Date: Mon Oct 22 00:03:57 2001, Shift: 16:00-24:00
    Supervisor: DO, with: KH until 19:00
    Log book#10 pages 43 - 45
    
    Shift Summary:
    ------------------------------------------------------------
    1.Machine Status:
      Long and nice store from 18:20 to   
      MCR needed 1h or so to steer the beams
    
    2.Achievements:
             Trig1   Trig5  Trig 6  FS        MRS
      Runs 
      5465  257762  247789  143011  4d. 1/3B  35d. 4A 
      5466   56011   53054   30735  4d. 0     0 
      5467  246357  237396   77205  4d. 1/3B  35d. 4A
    
    New setting
      5468  231835  222537   72271  4d. 1/2B  35d. 4B
       STILL RUNNING   
    
    3.Comments - Problems: 
      No detector problems, they worked nicely during
      all the shift. 
      BBL 37 HV was 0 before data taking.
      DAQ IO and Run servers had to be stopped and restarted 
      to start run 5465. 
    
    Note: although the goal with field setting at 1/3 for FS
    and pol A for MRS was reached, I decided (with KH) to rerun
    for 75k trig6 more after the 0 field run because MCR called
    in the mean time to tell us that  they ended the steer just
    when run 5465 was stopped. Run 5467 should contain some 
    clean events (well, don't know if the steering has some 
    other effects than just influencing the rates...) 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Oct 22 2001 - 00:04:35 EDT